# Bayesian Optimization for Non-Convex Two-Stage Stochastic Optimization Problems

### v1

@@ 3.5 The two-step incumbent method
20:  Fit MAP hyperparameters of GP prior, $f_{\text{step-2}}\sim\mathcal{GP}$, on functions ${\mathbb{X}}\times{\mathbb{U}}\to\mathbb{R}$ as described in Section E.3
@@
22:  Optimize ${\bm{x}}^{*n_{\mathrm{tot},2}}\in\operatorname*{arg\,max}_{\bm{x}}\frac{1}{N_{u,\mathrm{rec}}}\sum_{{\bm{u}}\in{\mathbb{U}}_{\mathrm{MC},\text{rec}}}\mu_{\text{step-2}}^{n_{\mathrm{tot},2}}({\bm{x}},{\bm{u}})$
@@ 4 Experiments
To demonstrate the value of jointly optimizing the fixed design and adjustable variables, experiments are run on a variety of synthetic and real-world examples. Using the synthetic test problems sampled from Gaussian processes we avoid any issues due to model mismatch, investigate the effects of dimension and length scales, and demonstrate the ability of the algorithms to cope with observation noise. The example of optical table design shows that the joint and alternating knowledge gradient remain superior algorithms in real-world optimization landscapes.
In all experiments, the initial sample is a scrambled Sobol’ sequence, whose size depends on the problem dimension and is detailed separately in Sections 4.2 and 4.3.
The main algorithm in this paper, presented in Equation 12, jointly optimizes the fixed design ${\bm{x}}$ and control strategy ${\bm{g}}$ and we refer to it as joint KG (jKG). We refer to the alternating alternative from Section 3.4 as alternating KG (aKG) and the two-step incumbent from Section 3.5 as two-step KG (2sKG). In addition to comparing the three knowledge gradient algorithms, we also compare with a joint and a two-step random sampling policy, which we call joint RS (jRS) and two-step RS (2sRS) respectively. These policies use the same models and recommendation procedures as the joint and two-step knowledge gradient but recommend samples according to a scrambled Sobol’ sequence.
@@ 4.1 Performance Metric
$r^{n}\approx\hat{r}^{n}=\frac{1}{MN_{u,\text{rec}}}\sum_{i=1}^{M}\sum_{j=1}^{N_{u,\text{rec}}}\biggl{(}h({\bm{x}}^{*n,i},{\bm{g}}^{*n,i}({\bm{u}}^{j}),{\bm{u}}^{j})-h({\bm{x}}^{*},{\bm{g}}^{*}({\bm{u}}^{j}),{\bm{u}}^{j})\biggr{)}.$ (21)
@@
In the results that follow, we plot this estimate as a function of $n$ for each algorithm, along with a 95% confidence interval calculated as two standard errors on either side of the mean.
@@ 4.2 Synthetic examples: random GP samples
The synthetic examples are generated by sampling a Gaussian process (GP). By using GP samples, we remove model-mismatch problems from the picture, such as landscapes with heterogeneous length scales. For each family of problems, characterized by dimension and length scale, 100 different test problems are generated by taking 100 independent samples of the GP. The samples are generated by randomly sampling the weights and basis vectors for 1024 random Fourier features as in (Rahimi & Recht, 2007), using the implementation in BoTorch (Balandat et al., 2020).
@@
Figure 2: Evolution of the expected simple regret for GP sampled problems of different dimensions, with gray shading showing the evaluations attributed to the initial design. Note that the joint and alternating algorithms only use the first initial design phase, while the two-step algorithms use both. Joint KG and alternating KG perform well in all cases, with joint KG approaching zero regret slightly faster than alternating KG. Both algorithms outperform the two-step and random sampling benchmark algorithms. The two-step algorithms do particularly badly when the $({\bm{x}},{\bm{y}},{\bm{u}})$-dimension is $(4,1,1)$ since the first step only optimizes the adjustable variables while the bulk of the optimization problem is in the fixed design space.
@@
Figure 3: Evolution of the expected simple regret for GP sampled problems of different length scales, with gray shading showing the evaluations attributed to the initial design. Note that the joint and alternating algorithms only use the first initial design phase, while the two-step algorithms use both. Joint KG and alternating KG are consistently the best performers, and the two-step algorithms perform particularly poorly when the short length scale is in the ${\bm{x}}$-dimension.
@@
Figure 4: Evolution of the expected simple regret for GP sampled problems with additive Gaussian observation noise of standard deviation $\sigma=2$. The problems have $({\bm{x}},{\bm{y}},{\bm{u}})$-dimension $(2,2,2)$ and come from the same distribution as those which appear in Figure 2. Gray shading shows the evaluations attributed to the initial design. Note that the joint and alternating algorithms only use the first initial design phase, while the two-step algorithms use both. While the performance of all algorithms gets worse with increased levels of observation noise, the joint and alternating KG continue to outperform the Sobol’ sequence and two-step benchmarks.
4.3 Real world example: design of an optical table
The optimization landscapes found in real-world problems are typically unlikely to be seen as samples from a stationary GP. To confirm that the joint and alternating KG continue to be the best performers in the presence of model-mismatch, we consider the design of an optical table. This example is a simplification of the example in (Salomon, 2019). An optical table is used in optics experiments to reduce the amplitude of vibrations propagated from the environment to the equipment.
@@ 4.3 Real world example: design of an optical table
For the stochastic optimization problem, the spring constants $k$ are taken as the fixed design, the vibration frequency of the floor, $\omega$, is the uncertain environmental variable, and the damping coefficient, $c$, is the adjustable variable. For the GP modeling, we approximate the negative logarithm of the amplitude ratio, $-\log_{10}(B/A)$, in place of the amplitude ratio as a function of $k$, $c$ and $\log_{10}(\omega)$.
Figure 5(b) shows that joint KG and alternating KG approach zero much faster than the Sobol’ and two-step algorithms. The two-step KG is initially the fastest decreasing because, in the first step, it is solving the simpler problem of finding the best damping coefficient, $c$, for a fixed spring constant, $k$. However, because it is not known a priori for how long to run each step, and because a new set of initial samples must be collected each time the algorithm switches steps, it cannot maintain this fast convergence rate.
@@
Figure 5: Schematic diagram and evolution of the simple regret for the optical table experiment. The joint and alternating KG algorithms are the fastest to converge to zero regret.
@@ 4.4 The cost of optimizing the acquisition function
It is important that the time required to optimize the acquisition function is negligible compared with the time or cost required to evaluate the expensive objective. As an example, on a (2,2,2)-dimensional test problem, running with 6 CPU cores on an Intel Xeon Platinum 826 2.9GHz processor it took an average of just 21 seconds to optimize the joint KG acquisition function.
Figure 6 shows the mean wall clock time for a single optimization of the acquisition function for each of the three knowledge gradient variants on the (2,2,2)-dimensional GP test problems. The alternating and two-step KG algorithms each involve two acquisition functions, and the figure shows the average of the time required to optimize each one.
Thanks to the Monte-Carlo approximation of the outer expectation in Equation 12, the joint KG algorithm is immediately parallelized when implemented in PyTorch. However, the exact expectation in the alternating (Equation 17 and Equation 15c) and two-step KG (Equation 19) require more work to be parallelized, which is beyond the scope of this paper.
Therefore, the timings in Figure 6 use 6 CPUs for the joint KG, but only 2 CPUs for the alternating and two-step KG since the implementations of the latter mean a larger number of CPUs do not improve the timings.
Figure 6: Mean wall times required to optimize the acquisition functions on the (2,2,2)-dimensional GP sampled test problems without observation noise. The joint KG is optimized using 6 CPU cores, while the remaining algorithms were given just 2 CPU cores because their implementation means they attain poor cluster utilization on higher numbers of cores.
@@ 5 Discussion
The results in Section 4 show the importance of jointly optimizing the fixed design and control strategy for the adjustable variables. The joint knowledge gradient is conceptually appealing and computationally feasible, while the alternating knowledge gradient provides a natural alternative and demonstrates that the Monte Carlo approximation of the expectation over the hypothesized next observation does not adversely affect the results. It is critically important in Bayesian optimization that optimizing the acquisition function is significantly cheaper than evaluating the expensive objective, $h$. The joint and alternating knowledge gradient algorithms are both applicable to settings where the expensive function takes several minutes or more to evaluate. If further efficiencies were required, then the computational strategy for the joint knowledge gradient would also be naturally suitable for running on a GPU.
There are many possible future research directions. Stochastic programming typically involves explicit constraints, and while the algorithm can support optimizing over constrained spaces ${\mathbb{X}}$ and ${\mathbb{Y}}$, it would be interesting to extend it to handle black-box constraints as in (Ungredda & Branke, 2024). Similarly, the algorithms could be extended to cover standard problems in stochastic programming such as multi-stage decisions and different risk measures such as value-at-risk.
Finally, another limitation of the algorithm is its extensibility to higher dimensions. For example, in the wind farm layout problem, there are typically hundreds of wind turbines leading to hundreds of parameters. High-dimensional Bayesian optimization is an active area of research (Binois & Wycoff, 2022) and recent ideas such as local search methods could be incorporated to tackle higher-dimensional two-stage stochastic programs.
@@ Acknowledgments
The authors would like to thank Andrew Nugent for his verification of the proofs in Appendix A and his suggestions for improving the clarity of our explanations.
@@ Appendix A Theoretical Results
In this section, we establish the asymptotic consistency of the value associated with the recommendations of the joint knowledge gradient acquisition function presented in Section 3 when viewed as an estimator for $\sup_{{\bm{x}},{\bm{g}}}\overline{h}({\bm{x}},{\bm{g}})$. This was stated in the main text as Section 3.1.
@@
The proof of consistency is based on the work by (Bect et al., 2019) and (Buckingham et al., 2023), but avoids the use of random measures which is found in (Bect et al., 2019). It begins by showing that the joint knowledge gradient converges almost surely, uniformly to zero in Theorem A.10. Further, the posterior mean and covariance of $\overline{f}$ are shown to converge uniformly to continuous limits in Proposition A.12 by observing that the posterior mean and covariance of $f$ are martingales in Banach spaces. These two facts are combined to show in Proposition A.16 that the uniform limit, $\overline{\mu}^{\infty}$, of the posterior means of $\overline{f}$ has sample paths which are almost surely equal to the sample paths of $\overline{f}$ (up to a constant). Whence, convergence of the values of the recommendations can be established in Theorem A.17.

### v2

@@ 3.4 The two-step incumbent method
20:  Fit MAP hyperparameters of GP prior, $f_{\text{step-2}}\sim\mathcal{GP}$, on functions ${\mathbb{X}}\times{\mathbb{U}}\to\mathbb{R}$
@@
22:  Optimize $\hat{{\bm{x}}}^{*n_{\mathrm{tot},2}}\in\operatorname*{arg\,max}_{\bm{x}}\frac{1}{N_{u,\mathrm{rec}}}\sum_{{\bm{u}}\in{\mathbb{U}}_{\mathrm{MC},\text{rec}}}\mu_{\text{step-2}}^{n_{\mathrm{tot},2}}({\bm{x}},{\bm{u}})$
@@ 4 Experiments
To demonstrate the value of jointly optimizing the fixed design and adjustable variables, experiments are run on a variety of synthetic and real-world examples. Using the synthetic test problems sampled from Gaussian processes we investigate the effects of dimension and length scales, and demonstrate the ability of the algorithms to cope with observation noise. This setting allows us to isolate these effects without potential interference from model mismatch. The examples of optical table design and supply-chain optimization validate the algorithms in real-world optimization landscapes, and allow us to compare to the current state of the art from (Xie et al., 2021).
In all experiments, the initial sample is a scrambled Sobol’ sequence, whose size depends on the problem dimension and is detailed separately in Section E.1.
The main algorithm in this paper, presented in Algorithms 2 and 12, jointly optimizes the fixed design ${\bm{x}}$ and control strategy ${\bm{g}}$ and we refer to it as joint KG (jKG). We refer to the alternating alternative from Section 3.3 as alternating KG (aKG) and the two-step incumbent from Section 3.4 as two-step KG (2sKG). In addition to comparing the three knowledge gradient algorithms, we also compare with a joint and a two-step random sampling policy, which we call joint RS (jRS) and two-step RS (2sRS) respectively. These policies use the same models and recommendation procedures as the joint and two-step knowledge gradient but recommend samples according to a scrambled Sobol’ sequence. The fixed design ${\bm{x}}_{\text{step-1}}$ used in the first step of the two-step algorithms is in the center of the space ${\mathbb{X}}$.
@@ 4.1 Performance metric
$\hat{r}^{n}=\frac{1}{MN_{u,\text{rec}}}\sum_{i=1}^{M}\sum_{j=1}^{N_{u,\text{rec}}}\biggl{(}h({\bm{x}}^{*},{\bm{g}}^{*}({\bm{u}}^{j}),{\bm{u}}^{j})-h({\bm{x}}^{*n,i},{\bm{g}}^{*n,i}({\bm{u}}^{j}),{\bm{u}}^{j})\biggr{)}.$ (21)
@@
In the results that follow, we plot this estimate as a function of $n$ for each algorithm and indicate the region of two standard errors on either side of the mean. By the Central Limit Theorem, this will approximate a 95% confidence interval.
@@ 4.2 Synthetic examples: random GP samples
The synthetic examples are generated by sampling a GP. By using GP samples, we remove model-mismatch issues from the picture, such as landscapes with heterogeneous length scales. For each family of problems, characterized by dimension and length scale, 100 different test problems are generated by taking 100 independent samples of the GP. The samples are generated by randomly sampling the weights and basis vectors for 1024 random Fourier features as in (Rahimi & Recht, 2007), using the implementation in BoTorch (Balandat et al., 2020). Note that while the GP test problems are sampled from a known distribution, we still fit the hyperparameters of the surrogate GP to the available observations at each iteration using MAP estimates, as described in Section E.4.
@@
Figure 2: Evolution of the expected simple regret for GP sampled problems of different dimensions, with gray shading showing the evaluations attributed to the initial design. Note that the joint and alternating algorithms only require the first initial design phase, while the two-step algorithms require both. The colored shaded regions around each line indicate two standard errors either side of the mean (an approximate 95% confidence interval). Joint KG and alternating KG perform well in all cases, with joint KG approaching zero regret slightly faster than alternating KG. Both algorithms outperform the two-step and random sampling benchmark algorithms. The two-step algorithms do particularly badly when the $({\bm{x}},{\bm{y}},{\bm{u}})$-dimension is $(4,1,1)$ since the first step only optimizes the adjustable variables while the bulk of the optimization problem is in the fixed design space.
@@
Figure 3: Evolution of the expected simple regret for GP sampled problems of different length scales, with gray shading showing the evaluations attributed to the initial design. Note that the joint and alternating algorithms only require the first initial design phase, while the two-step algorithms require both. The colored shaded regions around each line indicate two standard errors either side of the mean (an approximate 95% confidence interval). Joint KG and alternating KG are consistently the best performers, and the two-step algorithms perform particularly poorly when the short length scale is in the ${\bm{x}}$-dimension.
@@
Figure 4: Evolution of the expected simple regret for GP sampled problems with additive Gaussian observation noise of standard deviation $\sigma=2$. The problems have $({\bm{x}},{\bm{y}},{\bm{u}})$-dimension $(2,2,2)$ and come from the same distribution as those which appear in Figure 2. Gray shading shows the evaluations attributed to the initial design. Note that the joint and alternating algorithms only require the first initial design phase, while the two-step algorithms require both. The colored shaded regions around each line indicate two standard errors either side of the mean (an approximate 95% confidence interval). While the performance of all algorithms gets worse with increased levels of observation noise, the joint and alternating KG continue to outperform the Sobol’ sequence and two-step benchmarks.
4.3 Design of an optical table
The optimization landscapes found in real-world problems are unlikely to be seen as samples from a stationary GP. To confirm that the joint and alternating KG continue to be the best performers in the presence of model-mismatch, we consider the design of an optical table. This example is a simplification of the example in (Salomon, 2019). An optical table is used in optics experiments to reduce the amplitude of vibrations propagated from the environment to the equipment.
@@ 4.3 Design of an optical table
For the stochastic optimization problem, the spring constants $k$ are taken as the fixed design, the vibration frequency of the floor $\omega$ is the uncertain environmental variable, and the damping coefficient $c$ is the adjustable variable. For the GP modeling, we approximate the negative logarithm of the amplitude ratio, $-\log_{10}(B/A)$, in place of the amplitude ratio as a function of $k$, $c$ and $\log_{10}(\omega)$.
Figure 5(b) shows that joint KG and alternating KG approach zero much faster than the Sobol’ and two-step algorithms. The two-step KG is initially the fastest decreasing because, in the first step, it is solving the simpler problem of finding the best damping coefficient $c$ for a fixed spring constant $k$. However, because it is not known a priori for how long to run each step, and because a new set of initial samples must be collected each time the algorithm switches steps, it cannot maintain this fast convergence rate.
@@
Figure 5: Schematic diagram and evolution of the simple regret for the optical table experiment. Gray shading shows the evaluations attributed to the initial design. Note that the joint and alternating algorithms only require the first initial design phase, while the two-step algorithms require both. The colored shaded regions around each line indicate two standard errors either side of the mean (an approximate 95% confidence interval). The joint and alternating KG algorithms are the fastest to converge to zero regret.
4.4 Supply chain optimization
Supply chain management problems are a common application of stochastic programs in operations research. In this example, taken from (Xie et al., 2021), we compare with the state-of-the-art algorithm proposed in that work.
The supply chain in question concerns a simplified bio-pharmaceutical process in which the production of a clinical product consumes raw materials, including soy and a raw chemical material. Orders for soy have a long lead time and must be made before uncertainty on the week-to-week demand is resolved. The lead time for the raw chemical is much shorter and can be made after we have a good prediction of the demand. The initial soy order $x$ is the fixed design, the daily production quantity $y_{1}$ along with the $(s,S)$ ordering policy for the raw chemical form the adjustable variables ${\bm{y}}$ and the demand for the next four weeks are the uncertain environmental variables ${\mathbf{u}}$. The objective is to minimize the costs that arise from the initial soy order, orders of the raw chemical product, the storage cost of surplus product, and the large cost to subcontract to fulfill unmet demand. These are 10 per unit of soy, 5 per unit of raw chemical, 5 per unit of manufactured product stored until the next week, and 100 per unit of manufactured product which must be subcontracted. It takes a single unit of soy and a single unit of the raw chemical to make one unit of the manufactured product. Table 1 lists the different variables in the supply chain problem and Algorithm 5 details the simulation.
The problem is constrained so that the daily production cannot exceed $1/20$ of the soy quantity. Further, the $(s,S)$ ordering policy for the raw chemical is constrained such that $S>s$. When modeling the objective with a GP, we model $(s,S-s)$ rather than $(s,S)$ to convert the constraint $s<S$ into a box constraint. The fixed design and adjustable variables are further constrained to the finite sets specified in Table 1. To handle the constraint $y_{1}\leq x/20$, the acquisition function is optimized using multi-start constrained SLSQP, and the result is then rounded to the nearest feasible value. As in (Xie et al., 2021), when generating a recommended fixed design $x^{*n}$ as well as when evaluating the recommended control policy ${\bm{g}}^{*n}$, an exhaustive search is performed over all possible $x$ and ${\bm{y}}=(y_{1},s,S)$. The marginal distribution for the environmental variables in the initial sample is normal instead of uniform to reflect that ${\mathbf{u}}$ has a normal distribution. Full details are provided in Section E.3.
Table 1: Variables used in the supply chain management problem
Variable Symbol Type Feasible values / Distribution
Initial soy $x$ fixed design $x\in\{0,20,\dots,5000\}$
Daily production $y_{1}$ adjustable $y_{1}\in[0,x/20]\cap\mathbb{Z}$
Raw chemical ordering policy $(s,S)$ adjustable $s,S\in\{100,200,300,400,500\}$ s.t. $s<S$
Weekly demand $u_{1},u_{2},u_{3},u_{4}$ environmental $u_{1},\dots,u_{4}\sim\mathcal{N}(150,10^{2})$
Algorithm 5 Supply chain management simulation
0:  Soy order quantity $x$, target daily production level $y_{1}$, raw chemical ordering policy parameters $(s,S)$, weekly demand $u_{1},u_{2},u_{3},u_{4}$
1:  // Other variables: raw chemical inventory $r$, clinical product inventory $w$, cost objective $C$
2:  $r\leftarrow 100$, $C\leftarrow 10x$, $w\leftarrow 0$
3:  for $i=1,2,3,4$ do
4:     // For each week
5:     for $j=1,2,3,4,5$ do
6:        // For each working day
7:        // Top up the raw chemical inventory
8:        if $r<s$ then
9:           $C\leftarrow C+5(S-r)$
10:           $r\leftarrow S$
11:        end if
12:        // Produce clinical product
13:        $w_{\text{new}}\leftarrow\min(y_{1},x,r)$
14:        $x\leftarrow x-w_{\text{new}}$, $r\leftarrow r-w_{\text{new}}$, $w\leftarrow w+w_{\text{new}}$
15:     end for
16:     // Calculate costs for unmet demand / surplus product
17:     if $w\geq u_{i}$ then
18:        $w\leftarrow w-u_{i}$
19:        $C\leftarrow C+5*w$
20:     else
21:        $C\leftarrow C+100*(u_{i}-w)$
22:        $w\leftarrow 0$
23:     end if
24:  end for
24:  Cost, $C$
The objective is to minimize costs in the supply chain and the average cost over 100 repetitions is presented in Figure 6. Xie et al. (2021) only seek to recommend the best fixed design $x^{*n}$, and measure its performance using the best possible ${\bm{y}}$ for $x^{*n}$ (found by exhaustive search). For a fair comparison, the dotted lines show the same metric for our algorithms. The solid lines show our usual metric, which measures the combination of the recommended fixed design $x^{*n}$ and control policy ${\bm{g}}^{*n}$. The values for the algorithm of (Xie et al., 2021) are taken directly from their paper.
The joint knowledge gradient attains the best value found by (Xie et al., 2021) (after 2000 evaluations) after just 80 evaluations using the optimal control policy (their metric) and after 140 evaluations when using the recommended control policy. We note that even a random (Sobol’) sampling strategy beats the best value found in (Xie et al., 2021) after 480 evaluations which shows the benefit of modeling the whole design space with a single GP. In (Xie et al., 2021), expensive function evaluations made on the inner loop can only be used to inform the surrogate model used for that specific $x$, while a joint model can be informed by evaluations at all nearby locations.
Figure 6: Evolution of the supply chain costs over 100 repetitions of the experiment. Solid lines show the cost associated with the recommended fixed design ${\bm{x}}$ and control policy ${\bm{y}}={\bm{g}}({\bm{x}})$ as in the previous problems. Dotted lines show the cost associated with the recommended fixed design ${\bm{x}}$ if the optimal ${\bm{y}}$-values can be chosen later. This is the metric used in (Xie et al., 2021), which we include to facilitate a fair comparison. The right-hand figure shows a magnified view of the left-hand figure. Gray shading shows the evaluations attributed to the initial design. Note that the joint and alternating algorithms only require the first initial design phase, while the two-step algorithms require both. The colored shaded regions around each line indicate two standard errors either side of the mean (an approximate 95% confidence interval).
@@ 4.5 The cost of optimizing the acquisition function
The time required to optimize the acquisition function must be negligible compared with the time or cost required to evaluate the expensive objective. As an example, on a (2,2,2)-dimensional test problem, running with 6 CPU cores on an Intel Xeon Platinum 826 2.9GHz processor, the median time to optimize the joint KG acquisition function was just 13.5 seconds.
Figure 7 shows the median wall clock time for a single optimization of the acquisition function for each of the three knowledge gradient variants on the (2,2,2)-dimensional GP test problems. The average is taken over the 100 repeats and 350 iterations (or 300 iterations for the case of the two-step KG).
Thanks to the Monte-Carlo approximation of the outer expectation in Equation 12, the joint KG algorithm is immediately parallelized when implemented in PyTorch. However, the exact expectation in the alternating (Equations 17 and 15c) and two-step KG (Equation 19) require more work to be parallelized, which is beyond the scope of this paper.
Figure 7: Median wall times required to optimize the acquisition functions on the (2,2,2)-dimensional GP sampled test problems without observation noise using 6 CPU cores.
@@ 5 Discussion
The results in Section 4 show the importance of jointly optimizing the fixed design and control strategy for the adjustable variables. The joint knowledge gradient is simple, principled, and computationally feasible, with a theoretical guarantee of convergence and superior empirical results to the current state of the art (Xie et al., 2021). Furthermore, it provides not only the recommended fixed design but also a recommendation for the control policy for the adjustable variables. The alternating knowledge gradient provides a natural alternative and demonstrates that the Monte Carlo approximation of the expectation over the hypothesized next observation does not adversely affect the results.
There are many possible future research directions. Stochastic programming typically involves explicit constraints, and while the algorithm can support optimizing over constrained spaces ${\mathbb{X}}$ and ${\mathbb{Y}}$, it would be interesting to extend it to handle black-box constraints as in (Ungredda & Branke, 2024). Similarly, the algorithms could be extended to cover standard problems in stochastic programming such as multi-stage decisions, and different risk measures such as value-at-risk.
Finally, another limitation of the algorithm is its extensibility to higher dimensions. For example, in the wind farm layout problem (Chen et al., 2022), there are typically hundreds of wind turbines leading to hundreds of parameters. High-dimensional Bayesian optimization is an active area of research (Binois & Wycoff, 2022) and recent ideas such as local search methods could be incorporated to tackle higher-dimensional two-stage stochastic programs.
@@ Acknowledgments
The authors would like to thank Andrew Nugent for his verification of the proofs in Appendix A and his suggestions for improving the clarity of our explanations. We thank Wei Xie and Hua Zheng for providing their source code.
@@
Data Availability Statement
This work is entirely theoretical, there is no data underpinning this publication.
@@ Appendix A Theoretical Results
In this section, we establish the asymptotic consistency of the value associated with the recommendations of the joint knowledge gradient acquisition function presented in Section 3 when viewed as an estimator for $\sup_{{\bm{x}},{\bm{g}}}\bar{h}({\bm{x}},{\bm{g}})$. This was stated in the main text as Section 3.1.
@@
The proof of consistency is based on the work by (Bect et al., 2019) and (Buckingham et al., 2023), but avoids the use of random measures which is found in (Bect et al., 2019). It begins by showing that the joint knowledge gradient converges almost surely, uniformly to zero in Theorem A.10. Further, the posterior mean and covariance of $\bar{f}$ are shown to converge uniformly to continuous limits in Proposition A.12 by observing that the posterior mean and covariance of $f$ are martingales in Banach spaces. These two facts are combined to show in Proposition A.16 that the uniform limit, $\bar{\mu}^{\infty}$, of the posterior means of $\bar{f}$ has sample paths which are almost surely equal to the sample paths of $\bar{f}$ (up to a constant). Whence, convergence of the values of the recommendations can be established in Theorem A.17.